# How Does Quantization Affect Multilingual LLMs?

### v1

@@ 4.7 Human Evaluation
fr es ja ko Avg Ltn/IE $\neg$
W8 -7.4% 0.6% 7.4% -12.0% -2.8% -3.4% -2.3%
W8A8-sq -9.4% -7.4% -2.0% 4.0% -3.7% -8.4% 1.0%
Internal W4-g -16.6% -4.6% -16.0% -4.6% -10.5% -10.6% -10.3%
W8 0.6% -5.4% 12.0% 0.0% 1.8% -2.4% 6.0%
W8A8-sq -7.4% -8.6% 0.0% -3.4% -4.8% -8.0% -1.7%
Dolly W4-g -9.4% -1.4% 2.6% -8.0% -4.1% -5.4% -2.7%
@@ 5 Related Work
There is a scarcity of research examining the impact of compression and quantization on multilingual tasks. Paglieri et al. (2024) study the impact of multilingual calibration sets on quantization, but their evaluation is English-only. Ramesh et al. (2023) study the effect of compression on multilingual model fairness in terms of classification accuracy, showing that while monolingual evaluation indicates a negative impact, multilingual evaluation differs across languages and dimensions. Kharazmi et al. (2023) show that recovering compression-caused performance loss of LSTMs is harder in a multilingual setting than monolingually. In machine translation, Diddee et al. (2022) show that distillation has a varied effect by language due to dependence on priors such as amount of synthetic data used and confidence of the teacher models, while quantization exhibits more consistent performance trends across languages. Our work is the first, to our knowledge, to study the effect of quantization on LLMs and for open-ended generation.
More broadly, multilingual data is an example of long tail data. Prior work shows that compression techniques like quantization and sparsity amplify disparate treatment of long-tail rare features (Hooker et al., 2019; Ahia et al., 2021; Ogueji et al., 2022; Hooker et al., 2020). Ogueji et al. (2022) show that depending on how out of distribution the task data is, sparsity-based compression can sometimes avoid overfitting to the training data, making a model better suited to the downstream task. Ahia et al. (2021) find that sparsity preserves machine translation performance on frequent sentences, but disparately impacts infrequent sentences.
@@
A recent line of work has emerged on techniques to improve performance of quantized LLMs, with the sole focus on English models and data for tuning and evaluation (Ahmadian et al., 2024; Dettmers et al., 2022; Xiao et al., 2023; Bondarenko et al., 2024; Gong et al., 2024). Even the most recent (Li et al., 2024; Liu et al., 2024) omit the multilingual dimension without acknowledging the limitation. Multilinguality and compression are both integral parts of LLMs, and our work explores this new territory.
@@
We consider how design choices such as quantization impact performance for users of different languages. A wider body of work examines how design choices impact performance on underrepresented features or subgroups. Zhuang et al. (2021) and Nelaturu et al. (2023) find that hardware choice incurs disparate impact on underrepresented features. Wang et al. (2022) establish that distillation imposes similar trade-offs, but the disproportionate harm to the long-tail could be mitigated by modifying the student-teacher objective. Ko et al. (2023) evaluate the positive role of ensembling disproportionately favoring underrepresented attributes. Bagdasaryan and Shmatikov (2019) show that differential privacy techniques like gradient clipping and noise injection disproportionately impact underrepresented features.
@@ 6 Conclusion & Future Work
We examine widely adopted quantization techniques for model compression and ask, How does quantization impact different languages? We perform an extensive study of quantization in state-of-the-art multilingual LLMs—from 8 billion to 103 billion parameters—in 20+ languages using automatic metrics, LLM-as-a-Judge, RM-as-a-Judge, and human evaluation. We find that: (1) Damage from quantization is much worse than appears from automatic metrics: even when not observed automatically, human evaluators notice it. (2) Quantization affects languages to varying degrees, with non-Latin script languages more severely affected on automatic benchmarks. (3) Challenging tasks degrade fast and severely: math performance is strikingly reduced, as are responses on realistic challenging prompts judged by humans. On a bright note, quantization occasionally brings performance benefits.
Our results urge attention to multilingual performance at all stages of system design. Researchers might extend our work to consider the impact of other decisions on multilingual performance, including on languages excluded from training and out-of-distribution tasks. By being mindful of the impact on long-tail features, we’ll build better systems to serve the world.
@@ 7 Limitations
Due to the number of methods, languages, and benchmarks we examine, we focus our evaluation on models from two families (Command R and Aya). As we observe similar trends across these models, our findings are likely to generalize to other LLMs. Nevertheless, models that have been optimized differently or trained with a focus on specific tasks such as code or mathematical reasoning may behave differently.
@@
For our study, we focused on languages that were supported by the models we evaluated. Performance deterioration is likely even larger for languages that are not or severely under-represented in the pre-training data. For such languages, evaluation is also more challenging due to poor availability of benchmark data and human annotators.
@@ A.2 Automatic Tasks - Full Results
de en es fr ja ru zh Avg Avg (-en) Ltn/IE $\neg$
W8 -11.7% -2.9% 4.8% 8.6% 7.0% -1.4% 8.7% 1.9% 2.7% -0.3% 4.8%
Aya-23-35b W4 -4.5% -3.5% -6.2% -1.4% -19.3% -7.6% -4.7% -6.8% -7.3% -3.9% -10.5%
W8 -2.0% -5.0% 0.9% 0.0% 6.2% 2.1% 9.8% 1.7% 2.8% -1.5% 6.0%
Aya-23-8b W4 -2.0% -11.7% -7.1% -12.4% -43.8% -11.6% 9.8% -11.2% -11.2% -8.3% -15.2%
Table A6: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on MGSM (5-shot). Ltn/IE are Latin-script/Indo-European languages: de, en, es, fr. $\neg$ are the rest: ja, ru, zh.
Avg (-en) Avg ar cs de el en es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh
35b FP16 77.6 77.9 78.9 78.2 77.1 76.4 84.7 81.0 75.8 81.9 65.6 77.8 79.8 75.9 73.3 77.7 75.8 83.8 78.9 79.6 74.1 77.6 78.3 81.2
W8 77.1 77.4 77.3 78.8 77.2 76.6 84.6 80.8 74.9 82.4 65.6 77.6 80.8 74.8 73.7 77.6 74.8 82.9 77.1 78.9 72.0 77.2 77.0 80.3
W4 73.0 73.4 73.8 74.9 73.2 70.8 83.2 77.0 71.4 78.1 61.0 73.9 76.2 71.7 67.4 73.0 70.4 80.1 74.3 73.3 68.2 73.0 71.4 78.8
8b FP16 64.8 65.3 65.6 61.9 65.6 64.0 77.0 67.0 63.6 69.6 54.3 67.4 65.7 65.2 61.7 63.8 61.3 69.1 65.7 69.7 58.1 66.8 62.3 72.2
W8 64.6 65.1 64.3 61.8 64.8 63.0 76.1 67.4 63.9 70.4 54.2 67.4 64.6 65.4 61.4 64.3 59.8 68.7 65.4 68.7 58.1 67.0 63.7 71.8
W4 59.3 59.9 61.9 57.0 61.6 57.7 73.8 61.1 58.2 65.7 49.8 64.7 58.3 60.7 51.1 60.7 54.9 62.0 59.8 63.9 50.1 61.0 58.8 66.2
@@
ar cs de el en es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh Avg Avg (-en) Ltn $\neg$
W8 -2.0% 0.7% 0.1% 0.2% -0.1% -0.3% -1.2% 0.7% 0.0% -0.3% 1.3% -1.5% 0.5% -0.1% -1.3% -1.1% -2.3% -0.8% -2.8% -0.4% -1.7% -1.1% -0.6% -0.6% -0.6% -0.7%
35b W4 -6.5% -4.3% -5.0% -7.4% -1.7% -4.9% -5.7% -4.6% -7.0% -5.0% -4.5% -5.6% -8.0% -6.0% -7.0% -4.4% -5.8% -7.8% -7.9% -5.9% -8.8% -3.0% -5.8% -6.0% -5.4% -6.3%
W8 -1.9% -0.2% -1.2% -1.6% -1.2% 0.7% 0.5% 1.3% -0.2% 0.0% -1.7% 0.3% -0.4% 0.9% -2.5% -0.6% -0.4% -1.4% 0.0% 0.3% 2.1% -0.6% -0.3% -0.3% -0.2% -0.5%
8b W4 -5.6% -7.9% -6.1% -9.9% -4.2% -8.8% -8.4% -5.6% -8.4% -4.1% -11.2% -7.0% -17.1% -4.9% -10.5% -10.3% -9.0% -8.3% -13.8% -8.7% -5.7% -8.3% -8.3% -8.5% -7.8% -9.1%
Table A8: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on Belebele. Ltn are Latin-script languages: cs, de, es, en, es, fr, id, it, nl, pl, pt, ro, tr, vi. $\neg$ are the rest.
ar de es fr it ja ko pt zh Avg
103B FP16 64.0 68.3 68.7 68.0 69.3 64.4 62.3 70.0 65.0 66.7
W8 64.1 68.3 68.7 68.1 69.4 64.3 62.3 69.9 65.0 66.7
W8A8-sq 63.5 67.9 68.8 68.0 69.1 63.6 61.8 69.2 64.9 66.3
W8A8 62.6 67.1 68.2 67.4 68.3 62.9 60.8 68.7 64.1 65.6
W4-g 62.9 67.5 68.2 67.6 68.6 62.8 61.1 68.6 64.0 65.7
W4 60.5 65.7 66.5 65.4 66.6 61.1 59.3 66.7 62.1 63.8
35B FP16 56.5 60.7 62.3 61.8 62.0 56.4 54.8 62.0 57.9 59.4
W8 56.5 60.6 62.2 61.8 61.9 56.4 54.7 62.1 57.9 59.3
W8A8 56.4 60.5 62.5 61.9 62.0 55.8 54.5 61.8 58.1 59.3
W4-g 55.4 59.7 62.0 61.0 60.7 54.4 53.2 60.8 56.6 58.2
@@
ar de es fr it ja ko pt zh Avg Ltn/IE $\neg$
W8 0.2% 0.0% 0.0% 0.1% 0.1% -0.2% 0.0% -0.1% 0.0% 0.0% 0.0% 0.0%
W8A8-sq -0.8% -0.6% 0.1% 0.1% -0.3% -1.3% -0.8% -1.1% -0.2% -0.5% -0.4% -0.8%
W8A8 -2.2% -1.8% -0.7% -1.0% -1.5% -2.3% -2.4% -1.8% -1.4% -1.7% -1.3% -2.1%
W4-g -1.7% -1.2% -0.7% -0.6% -1.0% -2.5% -1.9% -2.0% -1.5% -1.5% -1.1% -1.9%
103B W4 -5.5% -3.8% -3.1% -3.9% -3.8% -5.1% -4.8% -4.8% -4.4% -4.4% -3.9% -4.9%
W8 0.0% -0.2% -0.2% 0.0% -0.2% 0.0% -0.2% 0.2% 0.0% -0.1% -0.1% 0.0%
W8A8 -0.2% -0.3% 0.3% 0.2% 0.0% -1.1% -0.5% -0.3% 0.3% -0.2% 0.0% -0.4%
35B W4-g -1.9% -1.6% -0.5% -1.3% -2.1% -3.5% -2.9% -1.9% -2.2% -2.0% -1.5% -2.7%
Table A10: Relative performance (%$\Delta$) vs. FP16 for Command Models on mMMLU. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
Avg (-en) Avg ar de en es fr hi id it nl pt ro ru uk vi zh
FP16 58.2 58.8 53.9 60.4 66.7 61.6 62.0 47.8 58.9 61.5 60.3 62.0 59.7 57.8 56.3 55.3 57.5
Aya-23-35b W8 57.9 58.5 53.8 60.0 66.2 61.7 61.7 47.4 58.7 61.1 60.0 61.6 59.1 57.5 56.1 54.9 57.5
W4 56.6 57.2 52.3 58.7 65.2 60.3 60.4 45.7 57.4 59.8 58.6 60.5 57.7 56.5 55.0 53.8 56.1
FP16 48.2 48.6 45.1 50.0 54.6 50.9 51.0 39.7 48.8 50.7 49.7 50.8 49.9 47.8 46.8 46.5 47.1
Aya-23-8b W8 47.8 48.2 44.9 49.9 54.2 50.5 50.6 39.4 48.5 50.2 49.4 50.6 49.2 47.4 46.3 45.7 46.4
W4 46.7 47.1 43.9 48.4 53.4 49.4 49.0 38.4 47.5 49.1 47.9 49.1 48.0 46.2 45.6 44.9 46.1
@@
ar de en es fr hi id it nl pt ro ru uk vi zh Avg Avg (-en) Ltn $\neg$
W8 -0.2% -0.7% -0.6% 0.2% -0.4% -0.7% -0.3% -0.6% -0.5% -0.6% -1.1% -0.6% -0.3% -0.8% -0.1% -0.5% -0.5% -0.5% -0.4%
Aya-23-35b W4 -3.1% -2.8% -2.2% -2.0% -2.5% -4.4% -2.5% -2.7% -2.8% -2.5% -3.3% -2.2% -2.2% -2.7% -2.6% -2.7% -2.7% -2.6% -2.9%
W8 -0.5% -0.3% -0.8% -0.8% -0.8% -0.9% -0.6% -1.0% -0.7% -0.4% -1.4% -0.9% -1.1% -1.7% -1.5% -0.9% -0.9% -0.8% -1.0%
Aya-23-8b W4 -2.7% -3.3% -2.1% -2.9% -3.9% -3.4% -2.6% -3.2% -3.7% -3.3% -3.8% -3.4% -2.7% -3.5% -2.2% -3.1% -3.2% -3.2% -2.9%
Table A12: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on mMMLU (Okapi). Ltn are Latin-script languages: de, en, es, fr, id, it, nl, pt, ro, vi. $\neg$ are the rest.
@@
W4-g 2.8% -0.5% 5.4% 0.5% 1.9% -1.3% 1.7% 0.6% -2.2% 1.0% 1.6% 0.3%
Table A19: Relative performance (%$\Delta$) vs. FP16 for Command Models on Language Confusion metrics. Ltn/IE are Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.

### v2

@@ 4.7 Human Evaluation
non-English Stats
fr es ja ko en avg Ltn/IE $\neg$
W8 -7.4% 0.6% 7.4% -12.0% -4.0% -2.8% -3.4% -2.3%
W8A8-sq -9.4% -7.4% -2.0% 4.0% 6.6% -3.7% -8.4% 1.0%
Internal W4-g -16.6% -4.6% -16.0% -4.6% -7.4% -10.5% -10.6% -10.3%
W8 0.6% -5.4% 12.0% 0.0% -6.0% 1.8% -2.4% 6.0%
W8A8-sq -7.4% -8.6% 0.0% -3.4% 2.0% -4.8% -8.0% -1.7%
Dolly W4-g -9.4% -1.4% 2.6% -8.0% -10.0% -4.1% -5.4% -2.7%
@@ 5 Related Work
There is a scarcity of research examining the impact of compression and quantization on multilingual tasks. Paglieri et al. (2024) examine multilingual calibration sets, but their evaluation is English-only. Ramesh et al. (2023) study compression vis-a-vis multilingual fairness, showing that performance differs across languages and dimensions. Kharazmi et al. (2023) show that recovering compression-caused performance loss of LSTMs is harder multilingually than monolingually. In machine translation, distillation has varied effects by language related to priors such as amount of synthetic data used and confidence of the teacher models, while quantization exhibits more consistent trends across languages (Diddee et al., 2022). To our knowledge, ours is the first to study quantized LLMs for open-ended multilingual generation.
Multilingual data is an example of long tail data. Prior work shows that compression techniques like quantization and sparsity amplify disparate treatment of long-tail rare features (e.g. Hooker et al., 2019; Ahia et al., 2021; Ogueji et al., 2022; Hooker et al., 2020). Similar to our observation of occasional performance gain, Ogueji et al. (2022) show that sparsity-based compression sometimes makes a model better suited to the downstream task. Ahia et al. (2021) find that sparsity preserves machine translation performance on frequent sentences, but disparately impacts infrequent sentences. Badshah and Sajjad (2024) also report some performance gain at lower precision.
@@
Recent work to improve quantized LLMs solely focuses on English models and data for tuning and evaluation (e.g. Ahmadian et al., 2024; Dettmers et al., 2022; Xiao et al., 2023; Bondarenko et al., 2024; Gong et al., 2024). Dettmers and Zettlemoyer (2023) perform a fine-grained sweep across bit-widths (3-8 bit), data types and quantization methods, and recommended 4-bit as the optimal size-performance trade-off, but do not evaluate multilingually. Huang et al. (2024) extensively analyze quantized LLaMA3 models in English. Badshah and Sajjad (2024) examine the effect of 4-bit NormalFloat (Dettmers et al., 2023) and 8-bit LLM.int8() Dettmers et al. (2022) on across model sizes and a variety of English tasks, finding that larger models are more resilient to quantization and performs better than smaller models at higher precision. Even the most recent (e.g. Li et al., 2024; Liu et al., 2024) omit multilinguality without acknowledging the limitation.
@@
We consider how design choices like quantization impact performance for users of different languages. A wider body of work examines how design choices impact performance on underrepresented features or subgroups. Zhuang et al. (2021) and Nelaturu et al. (2023) find that hardware choice incurs disparate impact on underrepresented features. Wang et al. (2022) show that distillation imposes similar trade-offs, and that harm to the long-tail can be mitigated by modifying the student-teacher objective. Ko et al. (2023) show that ensembling disproportionately favors underrepresented attributes. Differential privacy techniques like gradient clipping and noise injection also disproportionately impact underrepresented features (Bagdasaryan and Shmatikov, 2019).
@@ 6 Conclusion & Future Work
We examine widely adopted quantization techniques for model compression and ask, How does quantization impact different languages? We perform an extensive study in state-of-the-art multilingual LLMs—from 8 billion to 103 billion parameters—in 20+ languages using automatic metrics, LLM/RM-as-a-Judge, and human evaluation. We find that: (1) Damage from quantization is much worse than appears from automatic metrics: even when not observed automatically, human evaluators notice it. (2) Quantization affects languages to varying degrees, with non-Latin script languages more severely affected on automatic benchmarks. (3) Challenging tasks degrade fast and severely (e.g. mathematical reasoning and responses to realistic challenging prompts). On a bright note, quantization occasionally brings performance benefits.
Our results urge attention to multilingual performance at all stages of system design and might be extended to consider, for instance, languages excluded from training and out-of-distribution tasks. By minding the impact on long-tail features, we’ll build better systems to serve the world.
@@ 7 Limitations
Due to the number of methods, languages, and benchmarks we examine, we focus our evaluation on models from two families (Command R/R+ and Aya 23). As we observe similar trends across these models, our findings are likely to generalize to other LLMs. Nevertheless, models that have been optimized differently or trained with a focus on specific tasks such as code or mathematical reasoning may behave differently.
@@
For our study, we focused on languages that were supported by the models we evaluated. Performance deterioration is likely even larger for languages that are not in the pre-training data, or are severely under-represented. For such languages, evaluation is also more challenging due to poor availability of benchmark data and human annotators.
@@ A.2 Automatic Tasks - Full Results
non-en Stats
de es fr ja ru zh en Avg Ltn/IE $\neg$
Aya-23-35b W8 -11.7% 4.8% 8.6% 7.0% -1.4% 8.7% -2.9% 2.7% 0.6% 4.8%
W4 -4.5% -6.2% -1.4% -19.3% -7.6% -4.7% -3.5% -7.3% -4.0% -10.5%
Aya-23-8b W8 -2.0% 0.9% 0.0% 6.2% 2.1% 9.8% -5.0% 2.8% -0.4% 6.0%
W4 -2.0% -7.1% -12.4% -43.8% -11.6% 9.8% -11.7% -11.2% -7.1% -15.2%
Table A6: Relative performance (%$\%\Delta$) vs. FP16 for Aya 23 models on MGSM (5-shot). Ltn/IE are non-English Latin-script/Indo-European languages: de, es, fr. $\neg$ are the rest: ja, ru, zh.
ar cs de el es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh en non-en Avg
35b FP16 78.9 78.2 77.1 76.4 81.0 75.8 81.9 65.6 77.8 79.8 75.9 73.3 77.7 75.8 83.8 78.9 79.6 74.1 77.6 78.3 81.2 84.7 77.6
W8 77.3 78.8 77.2 76.6 80.8 74.9 82.4 65.6 77.6 80.8 74.8 73.7 77.6 74.8 82.9 77.1 78.9 72.0 77.2 77.0 80.3 84.6 77.1
W4 73.8 74.9 73.2 70.8 77.0 71.4 78.1 61.0 73.9 76.2 71.7 67.4 73.0 70.4 80.1 74.3 73.3 68.2 73.0 71.4 78.8 83.2 73.0
8b FP16 65.6 61.9 65.6 64.0 67.0 63.6 69.6 54.3 67.4 65.7 65.2 61.7 63.8 61.3 69.1 65.7 69.7 58.1 66.8 62.3 72.2 77.0 64.8
W8 64.3 61.8 64.8 63.0 67.4 63.9 70.4 54.2 67.4 64.6 65.4 61.4 64.3 59.8 68.7 65.4 68.7 58.1 67.0 63.7 71.8 76.1 64.6
W4 61.9 57.0 61.6 57.7 61.1 58.2 65.7 49.8 64.7 58.3 60.7 51.1 60.7 54.9 62.0 59.8 63.9 50.1 61.0 58.8 66.2 73.8 59.3
@@
non-En Stats
ar cs de el es fa fr hi id it ja ko nl pl pt ro ru tr uk vi zh en Avg Ltn $\neg$
35b W8 -2.0% 0.7% 0.1% 0.2% -0.3% -1.2% 0.7% 0.0% -0.3% 1.3% -1.5% 0.5% -0.1% -1.3% -1.1% -2.3% -0.8% -2.8% -0.4% -1.7% -1.1% -0.1% -0.6% -0.6% -0.7%
W4 -6.5% -4.3% -5.0% -7.4% -4.9% -5.7% -4.6% -7.0% -5.0% -4.5% -5.6% -8.0% -6.0% -7.0% -4.4% -5.8% -7.8% -7.9% -5.9% -8.8% -3.0% -1.7% -6.0% -5.7% -6.3%
8b W8 -1.9% -0.2% -1.2% -1.6% 0.7% 0.5% 1.3% -0.2% 0.0% -1.7% 0.3% -0.4% 0.9% -2.5% -0.6% -0.4% -1.4% 0.0% 0.3% 2.1% -0.6% -1.2% -0.3% -0.1% -0.5%
W4 -5.6% -7.9% -6.1% -9.9% -8.8% -8.4% -5.6% -8.4% -4.1% -11.2% -7.0% -17.1% -4.9% -10.5% -10.3% -9.0% -8.3% -13.8% -8.7% -5.7% -8.3% -4.2% -8.5% -8.1% -9.1%
Table A8: Relative performance (%$\%\Delta$) vs. FP16 for Aya 23 models on Belebele. Ltn are non-English Latin-script languages: cs, de, es, es, fr, id, it, nl, pl, pt, ro, tr, vi. $\neg$ are the rest.
ar de es fr it ja ko pt zh en non-en Avg
103B FP16 64.0 68.3 68.7 68.0 69.3 64.4 62.3 70.0 65.0 75.7 66.7
W8 64.1 68.3 68.7 68.1 69.4 64.3 62.3 69.9 65.0 75.7 66.7
W8A8-sq 63.5 67.9 68.8 68.0 69.1 63.6 61.8 69.2 64.9 75.6 66.3
W8A8 62.6 67.1 68.2 67.4 68.3 62.9 60.8 68.7 64.1 75.2 65.6
W4-g 62.9 67.5 68.2 67.6 68.6 62.8 61.1 68.6 64.0 74.4 65.7
W4 60.5 65.7 66.5 65.4 66.6 61.1 59.3 66.7 62.1 73.2 63.8
35B FP16 56.5 60.7 62.3 61.8 62.0 56.4 54.8 62.0 57.9 67.7 59.4
W8 56.5 60.6 62.2 61.8 61.9 56.4 54.7 62.1 57.9 67.7 59.3
W8A8 56.4 60.5 62.5 61.9 62.0 55.8 54.5 61.8 58.1 67.7 59.3
W4-g 55.4 59.7 62.0 61.0 60.7 54.4 53.2 60.8 56.6 66.5 58.2
@@
non-en Stats
ar de es fr it ja ko pt zh en Avg Ltn/IE $\neg$
W8 0.2% 0.0% 0.0% 0.1% 0.1% -0.2% 0.0% -0.1% 0.0% 0.0% 0.0% 0.0% 0.0%
W8A8-sq -0.8% -0.6% 0.1% 0.1% -0.3% -1.3% -0.8% -1.1% -0.2% -0.1% -0.5% -0.4% -0.8%
W8A8 -2.2% -1.8% -0.7% -1.0% -1.5% -2.3% -2.4% -1.8% -1.4% -0.7% -1.7% -1.3% -2.1%
W4-g -1.7% -1.2% -0.7% -0.6% -1.0% -2.5% -1.9% -2.0% -1.5% -1.7% -1.5% -1.1% -1.9%
103B W4 -5.5% -3.8% -3.1% -3.9% -3.8% -5.1% -4.8% -4.8% -4.4% -3.3% -4.4% -3.9% -4.9%
W8 0.0% -0.2% -0.2% 0.0% -0.2% 0.0% -0.2% 0.2% 0.0% -0.1% -0.1% -0.1% 0.0%
W8A8 -0.2% -0.3% 0.3% 0.2% 0.0% -1.1% -0.5% -0.3% 0.3% 0.0% -0.2% 0.0% -0.4%
35B W4-g -1.9% -1.6% -0.5% -1.3% -2.1% -3.5% -2.9% -1.9% -2.2% -1.8% -2.0% -1.5% -2.7%
Table A10: Relative performance (%$\Delta$) vs. FP16 for Command Models on mMMLU. Ltn/IE are non-English Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.
ar de es fr hi id it nl pt ro ru uk vi zh en non-en Avg
FP16 53.9 60.4 61.6 62.0 47.8 58.9 61.5 60.3 62.0 59.7 57.8 56.3 55.3 57.5 66.7 58.2
Aya-23-35b W8 53.8 60.0 61.7 61.7 47.4 58.7 61.1 60.0 61.6 59.1 57.5 56.1 54.9 57.5 66.2 57.9
W4 52.3 58.7 60.3 60.4 45.7 57.4 59.8 58.6 60.5 57.7 56.5 55.0 53.8 56.1 65.2 56.6
FP16 45.1 50.0 50.9 51.0 39.7 48.8 50.7 49.7 50.8 49.9 47.8 46.8 46.5 47.1 54.6 48.2
Aya-23-8b W8 44.9 49.9 50.5 50.6 39.4 48.5 50.2 49.4 50.6 49.2 47.4 46.3 45.7 46.4 54.2 47.8
W4 43.9 48.4 49.4 49.0 38.4 47.5 49.1 47.9 49.1 48.0 46.2 45.6 44.9 46.1 53.4 46.7
@@
non-En Stats
ar de es fr hi id it nl pt ro ru uk vi zh en Avg Ltn $\neg$
Aya-23-35b W8 -0.2% -0.7% 0.2% -0.4% -0.7% -0.3% -0.6% -0.5% -0.6% -1.1% -0.6% -0.3% -0.8% -0.1% -0.6% -0.5% -0.5% -0.4%
W4 -3.1% -2.8% -2.0% -2.5% -4.4% -2.5% -2.7% -2.8% -2.5% -3.3% -2.2% -2.2% -2.7% -2.6% -2.2% -2.7% -2.6% -2.9%
Aya-23-8b W8 -0.5% -0.3% -0.8% -0.8% -0.9% -0.6% -1.0% -0.7% -0.4% -1.4% -0.9% -1.1% -1.7% -1.5% -0.8% -0.9% -0.8% -1.0%
W4 -2.7% -3.3% -2.9% -3.9% -3.4% -2.6% -3.2% -3.7% -3.3% -3.8% -3.4% -2.7% -3.5% -2.2% -2.1% -3.2% -3.4% -2.9%
Table A12: Relative performance (%$\%\Delta$) vs. FP16 for Aya Models on mMMLU (Okapi). Ltn are non-English Latin-script languages: de, es, fr, id, it, nl, pt, ro, vi. $\neg$ are the rest.
@@
Avg Rel. %Delta mMMLU MGSM Belebele
en Ltn All en Ltn All en Ltn All en Ltn All
W8 -1.2% -0.2% 0.5% -0.6% -0.5% -0.5% -2.9% 0.6% 2.7% -0.1% -0.6% -0.6%
Aya-23-35b W4 -2.5% -4.1% -5.3% -2.2% -2.6% -2.7% -3.5% -4.0% -7.3% -1.7% -5.7% -6.0%
W8 -2.3% -0.4% 0.5% -0.8% -0.8% -0.9% -5.0% -0.4% 2.8% -1.2% -0.1% -0.3%
Aya-23-8b W4 -6.0% -6.2% -7.6% -2.1% -3.4% -3.2% -11.7% -7.1% -11.2% -4.2% -8.1% -8.5%
Table A17: Relative performance of quantized Aya 23 models in English vs. other languages. All non-English languages (All), non-English Latin-script languages (Ltn).
@@
35B W4-g 2.8% -0.5% 5.4% 0.5% 1.9% -1.3% 1.7% 0.6% -2.2% 1.0% 1.6% 0.3%
Table A20: Relative performance (%$\Delta$) vs. FP16 for Command Models on Language Confusion metrics. Ltn/IE are non-English Latin-script/Indo-European languages: de, es, fr, it, pt. $\neg$ are the rest: ar, ja, ko, zh.